OpenStudy (anonymous):

Let (v_1,v_2,...,v_r) be a basis for an inner product space V. How can I show that the zero vector is the only vector in V that is orthogonal to all of the basis vectors?

OpenStudy (anonymous):

Hmmm, well, any particular vector is going to be a linear combination of the basis vectors.

OpenStudy (anonymous):

A vector is not going to be orthogonal to a basis vector which it comes from.

OpenStudy (anonymous):

However, the basis are going to be orthogonal to each other.

OpenStudy (anonymous):

So lets consider any vector \(\mathbf u\): \[ \mathbf u = \sum_{i=1}^rc_i\mathbf v_i \]If \(\mathbf u\neq \mathbf 0\), then that means there is some \(c_i\neq 0\).

OpenStudy (anonymous):

We will suppose that \(c_j\neq 0\).\[ \mathbf v_j \cdot \left(\sum_{i=1}^rc_i\mathbf v_i\right) = \mathbf v_j\cdot c_j\mathbf v_j \]Because for all \(i\neq j\) we get \(\mathbf v_j\cdot c_i\mathbf v_i = 0\).

OpenStudy (anonymous):

Based on the properties of dot product, we can say: \[ \mathbf v_j \cdot \mathbf u = c_j\|\mathbf v_j\|^2\neq 0 \]